OpenStudy (anonymous):

7 with the little 35 at the top * 7 with the little negative 21 at the top * 7 wiith the little negative 8 in the top right hand side of the 7. whats the answer?

OpenStudy (phi):

like this \[ 6^{35} \cdot 7^{-21} \cdot 7^{-8} \] ?

OpenStudy (anonymous):

yes and i need to be able to see the work if possible!

OpenStudy (phi):

Let me fix the typo \[ 7^{35} \cdot 7^{-21} \cdot 7^{-8} \]

OpenStudy (phi):

the number in the upper right is how many times you multiply 7 times itself for example \(7^3 \) means 7*7*7

OpenStudy (phi):

if the little number is negative, you divide by 7 that many times

OpenStudy (phi):

if you remember that 7/7 is 1 this question means figure out how many 7's are multiplied, and how many are divided.

OpenStudy (anonymous):

when you divide the negative and the positive will the answer become a negative or stay a positive?

OpenStudy (anonymous):

like when you divide the little number and the big number, Say when you divide 7 into negative 21 will the answer be a negative or a positive?